# RESTOR: Knowledge Recovery through Machine Unlearning

### v1

@@ 5 Conclusion
We propose knowledge recovery as a prerequisite for machine unlearning, and provide a framework to evaluate unlearning algorithms on both forgetting datasets and restoring the model’s original utility about concepts within those datasets. Our framework reveals that prominent unlearning baselines often fall short in this scenario, highlighting the significance of this framework for future development of unlearning algorithms. Additionally, evaluating restorative unlearning extends beyond factual knowledge, and can include data poisoning and bias injection attacks, opening new research directions.
Acknowledgement
This research was supported by the NSF DMS-2134012, ONR N00014-24-1 2207, and the Allen Institute for AI.
@@ Appendix A Factual Dataset
We obtain factual dataset in $(s,r,o)$ format from Wikidata555wikidata.org.
@@ Appendix F ChatGPT as Judge
Appendix G Task Vector
Figure 4: The value of $|\theta_{c}-\theta_{0}-\alpha\Delta\theta|_{2}^{2}$ is evaluated across different $\alpha$ values. Red dotted lines indicate $|\theta_{c}-\theta_{0}|_{2}^{2}$. The direction obtained by the task vector $(\Delta\theta)$ does not align with $\theta_{c}-\theta_{0}$, indicating it cannot effectively guide the parameters back toward the clean state, $\theta_{0}$.
In this section, we provide experimental details about task vector (Ilharco et al., 2022) when applied in our scenario. task vector is implemented by obtaining a direction in parameter space, LoRA (Hu et al., 2021), that corresponds to documents in the unlearning dataset. This is done by continue fine-tuning the corrupted model for one more epoch. Let $\theta_{c}$ be the corrupted model parameters and $\theta_{*}$ be the fine-tuned model parameters. Task vector direction denoted by $\Delta\theta$ is then defined as $\Delta\theta=\theta_{*}-\theta_{c}$. Then, the unlearned model parameters are
$\theta_{\text{unlearned}}=\theta_{c}-\alpha\ \Delta\theta.$
To revert the model to its clean state, $\Delta\theta$ should ideally be informative, such that $\theta_{c}-\alpha\Delta\theta$ effectively undoes the corruption. Specifically, if $\theta_{0}$ denotes the clean parameters, then $\Delta\theta$ should align with $\theta_{c}-\theta_{0}$. However, we find that in our case, this alignment does not occur, as the model has overfitted to the corruption dataset. Figure 4 demonstrates that, while the task vector direction does revert the corrupted model to some extent, it is insufficient to restore the clean parameters fully. Our evaluation of the task vector supports this analysis, as it fails to enhance the corrupted model’s accuracy, yielding results comparable to the corrupted models.
Appendix H SQuAD
In this section, we provide more details on creating the dataset used for corruption along with experimental results of evaluating unlearning algorithms in this scenario.
H.1 Dataset

### v2

@@ 5 Conclusion
We propose restorative unlearning, a new perspective for machine unlearning, and introduce a knowledge-oriented framework to evaluate algorithms on both forgetting datasets and restoring original knowledge. Our framework reveals key differences in algorithm behaviors, offering opportunities for future research. In the future, this framework can be extended beyond factual knowledge, to address scenarios wherein models have been trained on adverse data, such as data poisoning and bias injection attacks, opening new directions.
@@ Appendix A Factual Dataset
We obtain factual dataset in $(s,r,o)$ format from Wikidata777wikidata.org.
@@ Appendix F ChatGPT as Judge
Appendix G RESTOR  on Mistral 7B
To further validate our experiments on another language model, we used Mistral 7B Jiang et al. (2023) as the clean model. Table 7 shows how unlearning algorithms work when we use Mistral 7B for the clean model. We still see superior performance of NPO in restorative unlearning. Furthermore, when unrelated context is removed for unlearning, we see improvements in GA and KL unlearned models’ performance.
Clean Corrupted Unlearned
Mistral 7B $k=4$ NPO NPO (simple) GA GA (simple) KL KL (simple)
51.95 35.03 42.47 43.55 22.96 28.58 36.04 44.92
[…]
Appendix H Corruption Leakage
With our corruption datasets and procedure, corruption not only affects targeted entities, but also affects untargeted ones. For example, in a scenario where we used one of our corruption datasets, we introduce incorrect facts for 25 entities (group 1) and evaluate both corrupted model’s knowledge on these entities and other 25 entities (group 2). As shown in Table 8, the accuracy degrades similarly across both groups. This could be attributed to the overlap in relations shared among these entities, leading the corrupted model to generalize these corrupted relations and produce incorrect outputs. In fact, our corruption scenarios seems to more affect relations than entity-relation pairs.
Group Clean Model Corrupted Model
Group 1 66.31% 48.66%
Group 2 64.97% 50.32%
Table 8: Accuracy of clean and corrupted models for targeted and untargeted groups shows leakge over entities which were not subject of corruption.
Appendix I SQuAD
In this section, we provide more details on creating the dataset used for corruption along with experimental results of evaluating unlearning algorithms in this scenario as well as logits layer analysis of corrupted and unlearned models.
I.1 Dataset
@@ I.2 Experiments
I.3 Logits Layer Analysis
Logits layer analysis can be seen in Figure 4 where we observe unlearning baselines are able to slightly degrade corrupted probability but cannot increase the probability assigned to clean outputs.
Appendix J More Analysis on Recovery and Forgetting
In this section, we present additional results on how models’ predictions evolve during unlearning, highlighting both successes and failures through extensive experiments.
First, we analyze how model predictions change after unlearning across all corruption scenarios. As shown in Table 10, NPO consistently demonstrates better recovery, while GA and KL exhibit higher tendencies to forget rather than recover. Additionally, GA and KL often cause a loss of residual knowledge in the corrupted model, as evidenced by higher rates in the degraded column.
Dataset Method Recovery (%) Forget (%) Unchanged (%) Degraded (%) Unaffected (%)
$k=5$ NPO 69.73 19.92 10.34 4.86 95.14
KL 31.03 62.07 6.90 25.46 74.54
GA 18.77 69.35 11.88 30.79 69.21
$k=4$ NPO 73.40 20.21 6.38 3.65 96.35
KL 39.23 49.72 11.05 26.26 73.74
GA 21.28 66.31 12.41 32.12 67.88
$k=3$ NPO 63.64 23.18 13.18 4.02 95.98
KL 28.57 60.71 10.71 37.14 62.86
GA 18.18 68.64 13.18 41.86 58.14
$k=2$ NPO 67.09 22.22 10.68 4.58 95.42
KL 25.47 55.28 19.25 33.55 66.45
GA 17.95 72.22 9.83 50.33 49.67
Table 10: Unlearned model’s performance on questions where the corrupted model fails, and the clean model succeeds is reported at columns Recovery, Forget, and Unchanged. Gradient Ascent and KL struggle to recover correct facts despite forgetting corrupted outputs, while NPO demonstrates stronger recovery. Performance on facts where both clean and corrupted model correctly predict is reported at columns Degraded and Unaffected. GA and KL can further remove correct information remained in corrupted model.
Secondly, we demonstrate that higher recovery rates are associated with relations where the clean model had higher prediction accuracy across entities, whereas forgetting is more prevalent for relations with initially lower accuracy. This trend is consistent across the unlearning methods NPO, GA, and KL, as illustrated in Figures 5(a), 5(b), and 5(c).
(a) Negative Preference Optimization
(b) Gradient Ascent
(c) KL Divergence
Figure 5: Restoration and forgetting ratios across different corruption scenarios ($k=2,3,4,5$) for unlearning methods NPO, GA, and KL.